the total revenue function for the two goods is given by the equation; TR=36x+3x²+56y-4y². Find the number of units of each good which must be sold if profit is to be maximised when the firm is subject to a budget constraint, 5y+10y=80. I am look for the solution to this problem


1
Expert's answer
2021-06-23T09:34:07-0400

Given

"TR=36x+3x\u00b2+56y-4y\u00b2"

Budget constant "5y+10y=80"

"x+2y=16\\x=16-2y"

"TR=36(16-2y)+3(16-2y)\u00b2+56y-4y\u00b2"

"\frac{∆TR}{∆Y}=-72y+6(16-2y)+56+8y"

"\frac{∆TR}{∆Y}=-104y+248=0"

"Y=\frac{248}{104}"

"x=16-\frac{496}{104}=\frac{1168}{104}"

"y=2.385\\x=11.231"